What kind of conic section the following quadratic form represents and transform it to principal axes.

4x21 + 6x1x2 − 4x22 = 10.


1
Expert's answer
2021-06-07T11:53:12-0400

We have "Q=x^TAx," where


"A=\begin{bmatrix}
 4 & 3 \\
 3 & -4
\end{bmatrix}, x=\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}"

"\begin{vmatrix}
 4-\lambda & 3 \\
 3 & -4-\lambda
\end{vmatrix}=0"


"(4-\lambda)(-4-\lambda)-9=0"

"\lambda^2-25=0"

"\lambda_1=-5, \lambda_2=5"

These are the eigenvalues.

Hence


"Q=-5y_1^2+5y_2^2"

We see that "Q=10" represents the hyperbola "-5y_1^2+5y_2^2=10"

"\lambda=-5"


"\begin{bmatrix}
 4-\lambda & 3 \\
 3 & -4-\lambda
\end{bmatrix}=\begin{bmatrix}
 9 & 3 \\
 3 & 1
\end{bmatrix}"

"R_1=R_1\/3"

"\begin{bmatrix}
 1 & 1/3 \\
 3 & 1
\end{bmatrix}"

"R_2=R_2-3R_1"


"\begin{bmatrix}
 1 & 1/3 \\
 0 & 0
\end{bmatrix}"

If we take "x_2=t," then "x_1=-t\/3."

Thus

"x=\begin{bmatrix}
 -t/3\\
 t
\end{bmatrix}=\begin{bmatrix}
 -1/3\\
 1
\end{bmatrix}t"


"\lambda=5"

"\begin{bmatrix}
 4-\lambda & 3 \\
 3 & -4-\lambda
\end{bmatrix}=\begin{bmatrix}
 -1 & 3 \\
 3 & -9
\end{bmatrix}"

"R_1=-R_1"

"\begin{bmatrix}
 1 & -3 \\
 3 & -9
\end{bmatrix}"

"R_2=R_2-3R_1"

"\begin{bmatrix}
 1 & -3 \\
 0 & 0
\end{bmatrix}"

If we take "x_2=t," then "x_1=3t."

Thus

"x=\begin{bmatrix}
 3t\\
 t
\end{bmatrix}=\begin{bmatrix}
 3\\
 1
\end{bmatrix}t"



"\dfrac{3}{\sqrt{10}}\begin{bmatrix}
 -1/3\\
 1
\end{bmatrix}=\begin{bmatrix}
 -1/\sqrt{10}\\
 3/\sqrt{10}
\end{bmatrix}"


"\dfrac{1}{\sqrt{10}}\begin{bmatrix}
 3\\
 1
\end{bmatrix}=\begin{bmatrix}
 3/\sqrt{10}\\
 1/\sqrt{10}
\end{bmatrix}"

Hence


"x=\begin{bmatrix}
 -1/\sqrt{10} & 3/\sqrt{10} \\
 3/\sqrt{10} & 1/\sqrt{10}
\end{bmatrix}\begin{bmatrix}
 y_1\\
 y_2
\end{bmatrix}"

"x_1=-\dfrac{1}{\sqrt{10}}y_1+\dfrac{3}{\sqrt{10}}y_2"

"x_2=\dfrac{3}{\sqrt{10}}y_1+\dfrac{1}{\sqrt{10}}y_2"